An athlete reaches the finish line at a velocity of 33 m s−1. She then applies a minimum braking force of 240 N as she moves along the uphill section of track to help her come to a stop.

Calculate the minimum uphill length of track L that should be available for braking. You should ignore all frictional forces other than those applied by the athlete.

mass of sledge and athlete = 95 kg


1
Expert's answer
2021-10-18T17:41:32-0400

"ma=-F_{br}=-240" N

"mv=-F_{br}t+c"


for t=0:

"v_0=33\ m/s \implies c=v_0m"


"mL=-F_{br}t^2\/2+v_0mt"


"L(t)=-F_{br}t^2\/(2m)+v_0t"


"L'(t)=F_{br}t\/m+v_0=0"


"t=mv_0/F_{br}=95\cdot 33/240=13" s


"L_{min}=33\cdot 13-240\cdot 13^2/(2\cdot 95)=215.5" m